OpenStudy (anonymous):

The length of a rectangle is 6 meters more than its width. If the area is 40 square meters, what are the length and width?

OpenStudy (anonymous):

\[L = W + 6\]\[A = LW=40\]\[40 = \left( W + 6 \right)W\]solve for W for the width then plug to find L for length